You roll two dice (each with six sides). What is the probability that the sum of the numbers on the dice are?
a. 3
b. 4
c. 7
d. 11

Answer by ewatrrr(24785)

 
Hi
___1_2_3__4__5__6
1|_2_3_4__5__6__7
2|_3_4_5__6__7__8
3|_4_5_6__7__8__9
4|_5_6_7__8__9_10
5|_6_7_8__9_10_11
6|_7_8_9_10_11_12
Just count them...(36 ways of rolling the two dice)
a. P(sum = 3) = 2/36 = 1/18
b. P(sum = 4) = 3/36 = 1/12
c. 7
d. 11